OpenStudy (vishweshshrimali5):

Nothing horrible looking equations this time - Math Challenge - 3

OpenStudy (vishweshshrimali5):

Let n be a given non negative integer. Determine the number S(n) of solutions of the equation: \[\large{x+2y+2z = n}\] in non negative integers x,y, z

OpenStudy (vishweshshrimali5):

@ikram002p @Miracrown

OpenStudy (anonymous):

@Zale101

OpenStudy (vishweshshrimali5):

I will post the answer to this question tomorrow :D Till then best of luck

OpenStudy (vishweshshrimali5):

@hartnn @JungHyunRan

Miracrown (miracrown):

so let's see if we can find a pattern for this one for example if n = 0, how many different solutions can there be? so n can be any non-negative integer 0,1,2,3,4,5,... and based on what n is, we want to know how many possible solutions there are for x,y,z given that x,y,z are also non-negative integers so, for example if n = 0, then there would be exactly 1 solution x = 0, y = 0, and z = 0 if n = 1 there is also exactly 1 solution x = 1, y = 0, and z = 0 if y or z is not zero, then it would be too big so what if n = 2 do you see how many solutions we would have? can you think of any possibilities for x,y, and z if n = 2?

OpenStudy (vishweshshrimali5):

Well tried @Miracrown but we would be lost very soon if we used that method. I speak from experience :( Or, may be I did something wrong *shrug*

Miracrown (miracrown):

you did something wrong, lol :)

OpenStudy (vishweshshrimali5):

Hahaha may be ? Uhm.. I solved this question at night with my eyelids closed :P There is a strong possibility *_*

OpenStudy (vishweshshrimali5):

* I solved using your approach at midnight

Miracrown (miracrown):

good on you!

OpenStudy (vishweshshrimali5):

Haha :D

OpenStudy (vishweshshrimali5):

I am not asking you to calculate all possible answers. I want you to find out a general number of solutions depending on the value of n. That is why I asked for S(n). There *are* infinitely many solutions :)

OpenStudy (vishweshshrimali5):

Hey no need of deleting it. It was perfectly correct till that step :) But, does not match my solution. But, there are many possible solutions

ganeshie8 (ganeshie8):

I'm getting this : \[\large S(n) = \lfloor\dfrac{n}{2} \rfloor + 1 + \dfrac{ \lfloor\dfrac{n}{2} \rfloor ( \lfloor\dfrac{n}{2} \rfloor +1)}{2} \]

OpenStudy (vishweshshrimali5):

Okay I am not going to check that but your answer and mine answer (the one given in the source) don't match. But that doesn't mean that your answer is wrong. I will post the solution tomorrow and we would discuss it then.

ganeshie8 (ganeshie8):

hmm this is a trivial linear diophantine equation with constraints, so solutions can be represented in several ways but verifying them should be easy enough...

ganeshie8 (ganeshie8):

i guess a nice word problem can be cooked up from the constraints :)

OpenStudy (anonymous):